Jason is now three times as old as helston. three years from now jason will be six times as old as helston one year ago find their present ages

Answer by Earlsdon(6294)

Let J = Jason's present age and H = Helston's present age.
1) J = 3*H "Jason is three times as old as Helston"
2) J+3 = 6(H-1) "Three years from now Jason will be six times as old as Helston (was) one year ago"
Substitute J from equation 1) into equation 2) and simplify.
(3*H)+3 = 6H-6 Subtract 3H from both sides.
3 = 3H-6 Add 6 to both sides.
9 = 3H Divide both sides by 3.
3 = H and...
J = 3*H = 3*3 = 9
Jason is 9 and Helston is 3